Science Operations Announcements

Patch to commissioning release for GMOS-S Hamamatsu CCDs

October 15th, 2014
This patch release of the "commissioning" release of the Gemini IRAF data reduction package is being made available to assist in the reduction of the Hamamatsu data, pending the next stable release of the Gemini IRAF package. In addition, this release is to allow users to provide valuable feedback as to the status of the data reduction package for GMOS-S Hamamatsu data. Please provide any feedback, comments or questions etc., via the Gemini HelpDesk system as described in the notes below for the v113_COMM release.

GNIRS Temporarily Unavailable

September 29th, 2014

Due to ongoing problems with its cold heads, GNIRS is about to be removed from the telescope and warmed up for closer inspection. We should know by Friday October 10th whether this intervention has been successful. The instrument will not be available for observations during this period, and NIRI will be installed on the GNIRS telescope port instead (see the 2014B telescope schedule: http://www.gemini.edu/sciops/metrics/gn2014Boverview.html).

Gemini North ODB offline Sep 24

September 23rd, 2014

The Gemini North observing database will be offline on September 24 from 9:45 am - 12:00 pm HST while we perform network maintenance. During this time the following services will be unavailable:

  • Phase I Tool guide star probability queries
  • CADC program key validation (Gemini North programs only)
  • Observing Tool synchronizations, file transfers, and non-sidereal target queries (Gemini North programs only)

Please check this site for further announcements.

Gemini North ODB offline Sep 11-17

September 9th, 2014

UPDATE Sep 17 @ 4:00pm - the Gemini North Observing Database is back online. Please file a Help Desk ticket if you encounter any problems.

UPDATE Sep 17 @ 8:30am - the new hardware has been installed and we are working with the vendor to resolve network issues. We are hopeful that the GN ODB will be back online later today.

UPDATE Sep 16 @ 8:15am - work is ongoing, and we are still offline. Will post an update when we have more details.

Pages